Grant may get a bonus next year of $150,000 if he achieves the greatest number of house sales in his employer's real estate business. If he comes second he will only get a bonus of $50,000 and if he doesn't come in the top two, he gets nothing. There is a 40% chance Grant gets first place, a 40% chance he gets second and a 20% chance he ranks below second place. What is the standard deviation of the bonus? (Round to the nearest dollar). Select one:

A. $48,166

B. $60,000

C. $80,000

D. $75,000
Answer #1

Expected bonus=40%*150000+40%*50000+20%*0=80000.00000

Standard deviation=sqrt(40%*(150000-80000)^2+40%*(50000-80000)^2+20%*(0-80000)^2)=60000.00000